Subject: 3.5-m Users Committee minutes 9/17/07

From: Bruce Gillespie

Submitted: Sun, 23 Sep 2007 18:25:14 -0600

Message number: 1058 (previous: 1057, next: 1059 up: Index)

                       APO 3.5-m Users Committee Phonecon, 9/17/07

Attending:  Suzanne Hawley, Bruce Gillespie, Mark Klaene, Scott  
Anderson, Remy Indebetouw, Ed Turner (for Michael Strauss), Russet  
McMillan, John Bally, Bill Ketzeback

Absent: Jon Holtzman, Al Harper, Jon Fulbright

Minutes taken by Bruce Gillespie


User feedback, comments from institutional representatives:

Princeton (Turner) - Ed said that Michael Strauss told him he had no  
user comments to share.  Since much of the time since the last  
meeting was during the summer shutdown, we would naturally expect  
there to be little user feedback.

UVa (Indebetouw) - Remy had nothing to report, other than UVa is  
organizing a site visit for their students this fall.  Russet and  
Suzanne mentioned that we will be hosting four separate student site  
visits at APO in the coming weeks.

UW (Anderson) - Scott had no user comments, and Suzanne mentioned  
that Paula Szkody thanked us for getting some guider images of a  
target that is scheduled for upcoming HST observations.

CU (Bally) - John had no user comments to report.

JHU (Fulbright) - No report

UChicago (Harper) - No report

NMSU (Holtzman) - No report


Discussion of telescope/instruments report - Klaene:

The report is given below; Mark walked through the highlights, and  
there was some discussion.

The summer shutdown started and finished on schedule, and the  
telescope and instruments are good-to-go for science observations.   
Nick McDonald, the recently hired 3.5-m telescope engineer, was given  
responsibility for most of the shutdown activities, with help from  
several others.  Nearly all the tasks were completed (see list  
below), and the leftovers will be attended to in the coming weeks.   
One of the small scallop cracks in the back of the primary mirror  
faceplate seems to have gotten somewhat larger, and as a precaution  
the mirror support at that cell has been disabled.  No discernible  
effect on the optical performance of the mirror has been detected.   
The PMSS was serviced and calibrated, and the instruments were all  
serviced except the echelle (which is serviced every two years).   
There may be a lingering problem with the DIS cryotigers that could  
require attention later this year, although they appear stable at the  
moment.  Except for NIC-FPS, all the instruments (including CorMASS  
and GFP) are up and ready for use.

Suzanne recently visited Kitt Peak while our mirror was there, and  
she mentioned that the mirror looked pretty good even before it was  
realuminized.  Mark said that the exact before/after improvement in  
reflectivity numbers is not yet available, but there was an  
indication that the reflectivity before recoating was ~91%, and only  
went up to ~92% afterward.  The scattered light from the surface did  
improve a factor of 10, from ~2% to about .2%.  This indicates that  
we are doing a pretty good job of keeping the mirror clean in the  
telescope for relatively long periods of time.  Russet said that some  
observers take calibrations off the scattered light from the primary  
mirror, and that they will need to increase their exposure times  

The DIS grating mounts were given new larger bushings to help protect  
the gratings during changes.  Also, the red-high grating has been  
shimmed to place the spectrum near the center of the detector.   
Suzanne asked if users had been having problems with the red-high  
spectrum placement in the past, and Russet said that some users had  
mentioned it during instrument checkout, but nobody had reported  
problems with data reductions.  With the spectrum now better  
positioned, the stray light from the on-chip amplifier glow is less  
of a problem.


   3.5-m Telescope, Instruments, and CIF Projects Highlights, 8/18/07  
through 9/12/07

                                   Mark Klaene

0) Overview

Monsoonal weather continued with moderate rain and evening clouds,  
which limited most of the on-sky time that would have been available  
for science observing. However, the shutdown started near the  
beginning of the reporting period, so little scheduled observing time  
was lost.  Wet weather continued through the shutdown, but we were  
largely able to work around it.  Construction started on the SDSS  
Support Building expansion project (for AS2) with no disruption to  
3.5-m work.

1) Telescope

The shutdown started on time and remained ahead of schedule for much  
of the period.  Weekend work was performed but limited in scope. To  
date all of the shutdown tasks (see list below) were completed except:

   o the Echelle fans have not been changed since the wrong parts  
were delivered, and

   o the new instrument rotator axis controller system was not tested  
on the telescope.

There were a few problems noted but none were particularly  
surprising.  The primary mirror cracks were inspected and were seen  
to have stayed basically unchanged except for the top of cell A3--the  
load cell for this support has been clamped off to relieve stress at  
this mirror position.  Loose flex pivots were noted on the tertiary  
mirror support and repaired.  Fifteen rivets were found to be loose  
and were replaced on the primary mirror covers.  The mirror covers  
were repainted and we anticipate higher overall cal lamp throughput  
as well as a noticeable improvement between 350 and 400 nm.  The new  
instrument rotator assembly was fit tested and aligned but could not  
go in to service due to continued work needed on the axis controller  

2) Instruments

   o The bushings were replaced on all DIS grating holders and the  
gratings were realigned.  This should reduce the risk of grating  
damage when installing gratings, especially in cold weather.  The red- 
high grating was moved into another mounting cell and a shim was made  
to position the spectrum on the middle of the chip;

   o Both DIS cameras were evacuated and the cryotiger lines  
cleaned.  The red cyrotiger compressor was replaced. Contamination  
was found in both sets of cryotiger lines;

   o SPIcam was evacuated and baked out;

   o NIC-FPS remains at CU for 16-channel readout and etalon  
upgrades.  We expect its return to the site in early October;

   o CorMASS was evacuated and cooled down; and

   o GFP was evacuated and cooled down.

                         Summer Shutdown Task List

M3:  Realuminize, service actuators

M1:  Realuminize, check and photograph cracks

PMSS:  Pressure transducer calibration, load cell calibration,  
bellofram inspection leak check, air line leak check with belloframs  
by sector, servo valves inspection/cleaning, clean mirror cell, vent  
tube repair, foam inspection, mirror cover actuator inspect and  
service, calibrate LVDTs, investigate fore/aft mirror constraint

Mirror covers:  Repaint covers, repair rivets

Pump SPIcam

Pump DIS:  Purge/recharge blue cam cryotiger system, install new  
bushings in grating holders, remount DIS red-high grating in new holder

Replace UPS batteries: Echelle, Inst - I level, Elect I -level, #3  
for ops bldg

Echelle electronics fan replacement (not completed)

New instrument rotator axis controller (not completed)

New NA2 rotator assembly

Repair strain gauge metrology noise issue (not completed)

Replace temperature sensors on primary mirror

Repair secondary mirror accelerometer recording system


Wide-field NIR instrument - Hawley, Gillespie:

Suzanne reported that we have received a second proposal from a non- 
ARC astronomer (the original proposal was made to us about 18 months  
ago) to build a wide-field NIR camera and grism spectrometer with a  
13-arcminute FOV and a programmable slit mask, and bring the  
instrument to the 3.5-m for their and our use.  This could be done in  
exchange for telescope time much like the arrangements we have used  
with CorMASS and the Goddard Fabry-Perot instruments; other  
collaborative scenarios are possible.  Several of the Users Committee  
members expressed strong support for continuing to pursue this  
initiative, and commented on how such an instrument would complement  
NIC-FPS and further our interests in general NIR science  
capabilities.  We need to determine a sense of how much ARC user  
interest there would be in having access to this instrument at APO  
before expending significant time or money on our side.  Users can  
check with their committee representatives for additional technical  
information about the instrument.

ACTION:  Users interested in a wide-field NIR camera at APO should  
send a paragraph or two to their Users Committee representative with  
copies to Suzanne and Bruce with an expression of interest and/or  
questions regarding this proposed instrument.


NIC-FPS status - Bally:

John reported that the NIC-FPS etalon has undergone multiple cryo  
cycles, and that the oscillations seem to be cured.  However, an  
accident happened two weeks ago involving the power supply and etalon  
motor controllers, which resulted in the need to replace the  
controllers, which will cause a delay in returning the instrument to  
APO.  Suzanne said that a new Q3 schedule, sans NIC-FPS, will come  
out shortly, and NIC-FPS will not be listed as available in Q4 until  
20 October.  The replacement controllers have been received in  
Boulder, and the instrument will be bench-tested with the 16-channel  
detector upgrade and etalon in a few days.  The science-grade and  
engineering-grade chips have been modified to allow the 16-channel  
reads, which will enable more efficient observations with Fowler  
sampling, and the electronics have been modified to reduce system  
noise.  Software work is underway both at CU and at APO to process  
the 16-channel reads, and to provide quick-look images during  
observing without excessive processing overheads.


Tertiary tip-tilt decision - Hawley:

Suzanne said that a decision was taken not to include tip-tilt  
capability into the new tertiary mirror support system, which is  
about to be designed.  The decision process used, which seemed to  
work and converge well, was to form a committee, solicit user input,  
examine existing tip-tilt systems on other telescopes, and look at  
existing telescope and site-intrinsic seeing data, to determine if  
tertiary tip-tilt should be a high priority for the 3.5-m telescope.   
Based on these evaluations, we will make fast guiding and auto-focus  
(using the existing NA2 guider) our near-term priorities.  If tip- 
tilt is determined to be needed in the future, we will employ the tip- 
tilt design hooks inherent in the new 2ndary mirror mounts.


Science Highlights from '07 - Hawley, Gillespie:

This is the time when we prepare the annual report for the ARC Board  
of Governors meeting.  One or two science highlights (text, graphics,  
images) of one or two pages for each highlight are traditionally  
presented in the report and orally to the Board members.

ACTION:  We need each institution's Users Committee member to poll  
their users and send these highlights to Suzanne and Bruce, not later  
than 15 October, for inclusion in the report.

For obvious reasons, we would like to have at least one  
representative user science highlight for each ARC institution.


Actions from earlier meetings - group

[open from previous months]:

ACTION:  Suzanne said that the Users Committee members need to  
identify which of their users would be willing to contribute effort  
and participate in a telecon to scope tip-tilt for the 3.5-m, and to  
communicate their names to her and Bruce Gillespie before 30 July.

STATUS:  Completed (see above), closed.

ACTION:  Jon Holtzman will investigate and report on the tip-tilt  
system at the TNG telescope on La Palma.

STATUS:  Completed (see above), closed.

[new actions from this month]:

ACTION:  Users interested in a wide-field NIR camera at APO should  
send a paragraph or two to their Users Committee representative with  
copies to Suzanne and Bruce with an expression of interest and/or  
questions regarding this proposed instrument.

ACTION:  We need each institution's Users Committee member to poll  
their users and send science highlights (since Oct. '06) to Suzanne  
and Bruce, not later than 15 October, for inclusion in the annual  
report to the ARC Board.


Next meeting:

The next Users Committee phonecon will be on Monday, 15 October, at  
8:30 AM Pacific Time.  The agenda and other materials will be sent to  
the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1058 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to